OpenStudy (anonymous):

What is the derivative of (-)tanhsech(5x+2)

OpenStudy (xapproachesinfinity):

what is that -tanhx(sexh(5x+2)

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

you wrote \(-tanh(sech(5x+2))\) it looked to me like this make sure!

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

so which one the first or the second? pls make sure

OpenStudy (xapproachesinfinity):

or do a snapshot of the problem so we are sure what we are doing

OpenStudy (anonymous):

the second one

OpenStudy (xapproachesinfinity):

So it is the composition of the composition

OpenStudy (xapproachesinfinity):

First let's work on the outer functions and then go to the inner functions okay using the Chain rule! do you know how to use Chain Rule?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

\(\frac{d}{dx}(f(g(x))=f'(g(x))g'(x)\)

OpenStudy (xapproachesinfinity):

let denote \(f(x)=-tanh(x),~~g(x)=sec(5x+2)\)

OpenStudy (xapproachesinfinity):

so we would have: \(\frac{d}{dx}(-tanh(sech(5x+2))=-tanh'(sech(5x+2))\times sech'(5x+2)\)

OpenStudy (xapproachesinfinity):

now what is (tanh(x))'?

OpenStudy (xapproachesinfinity):

Do you know it?

OpenStudy (anonymous):

sech^2x

OpenStudy (xapproachesinfinity):

correct! now instead of x we put sech(5x+2)

OpenStudy (xapproachesinfinity):

okay

OpenStudy (xapproachesinfinity):

??

OpenStudy (anonymous):

ok... thank you so much sir

OpenStudy (anonymous):

could you help with this If I = VR , and V = 250volts and R = 50 ohms. Find the chandes in I resulting from an increase of I volt in V and an increase of 0.5 0hms in R. I decreasesv by

OpenStudy (xapproachesinfinity):

now you have to use Chain rule one more time with (sech(5x+2))'

OpenStudy (xapproachesinfinity):

well that's easy find the initial current find the new one then find the difference like the volume problem

OpenStudy (anonymous):

ok sir

OpenStudy (anonymous):

what about this sir?

OpenStudy (anonymous):

Solve lim(x)→(π)(x2−sinx) . Find the limits

OpenStudy (xapproachesinfinity):

eh that's not clear! can try writing it one more time?

OpenStudy (anonymous):

limit as x tends to pi of (x^2 - sinx)

OpenStudy (xapproachesinfinity):

well that's easy no! sinx is defined for any number on the real line (real numbers) so just put pi into sin and x^2 (in fact any polynomial you just plug in the number to find the limit) because there are well defined every where

OpenStudy (xapproachesinfinity):

when the number that the the function is approaching is defined and is part of the domain then the limit is easy! just replace in the number you are trying to approach

OpenStudy (anonymous):

2π 1 π 0

OpenStudy (xapproachesinfinity):

again it is not clear

OpenStudy (anonymous):

2pi 1 pi 0

OpenStudy (xapproachesinfinity):

you sure! or make sure you wrote the function the right way

OpenStudy (anonymous):

am soooo sure

OpenStudy (xapproachesinfinity):

So let me ask this again this what you trying to find its limit \(\lim_{x\rightarrow\pi}(x^2-sin(x))\)

OpenStudy (xapproachesinfinity):

?

OpenStudy (anonymous):

very correct sir

OpenStudy (xapproachesinfinity):

in you option you should have \(\pi^2\)

OpenStudy (anonymous):

what i have is 2π 1 π 0

OpenStudy (anonymous):

2pi 1 pi 0

OpenStudy (xapproachesinfinity):

you know what post a snapshot of this problem okay!

OpenStudy (xapproachesinfinity):

something is not correct!

OpenStudy (xapproachesinfinity):

hey or you looking for limit of \(2x-sinx\) not \(x^2-sinx\)

OpenStudy (anonymous):

will, i do not hv the app sir

OpenStudy (xapproachesinfinity):

if you have the homework on you computer you can you use screen print button

OpenStudy (xapproachesinfinity):

print screen button*

OpenStudy (xapproachesinfinity):

or if it is a physical paper use the camera of your phone

OpenStudy (xapproachesinfinity):

hey you still there, pls be quick i have to go got bunch of homework to complete

OpenStudy (anonymous):

did you get it?

OpenStudy (xapproachesinfinity):

hey the options given are not correct! look at this graph and see what the function is approaching when x tends to pi https://www.desmos.com/screenshot/a1sw2fvzcc

OpenStudy (xapproachesinfinity):

pi=3,14..... is between 2 and 4

OpenStudy (xapproachesinfinity):

well btw 3 and 4 actually

OpenStudy (xapproachesinfinity):

should be \(\pi^2\)

OpenStudy (anonymous):

ok.... thanks so much. i hv to let you go do some other things..... will chat with you next time... God bless

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

bye friend and my tutor

OpenStudy (xapproachesinfinity):

you are welcome^_^ make sure to tell your instructor that there is a mistake there

OpenStudy (xapproachesinfinity):

bye

OpenStudy (anonymous):

yes sir